OpenStudy (anonymous):

I'm completely stumped on how to solve this one: 2\tan \left(\frac{x}{2}-\frac{\pi}{6} \right)= 5

OpenStudy (anonymous):

\[2\tan \left(\frac{x}{2}-\frac{\pi}{6} \right)= 5\]

OpenStudy (anonymous):

you need to simplify the fraction in the brackets

OpenStudy (anonymous):



OpenStudy (anonymous):

I'm presuming you mean \[\frac{x}{2} = \frac{3x}{6}\] instead of \[\frac{\pi}{2} = \frac{3\pi}{6}\]?

OpenStudy (anonymous):

oh sure i thought it was a pi. so yes

OpenStudy (anonymous):

actually in that case ....

OpenStudy (anonymous):